# DocNet: Semantic Structure in Inductive Bias Detection Models

### v1

@@ Abstract
News will have biases so long as people have opinions. However, as social media becomes the primary entry point for news and partisan gaps increase, it is increasingly important for informed citizens to be able to identify bias. People will be able to take action to avoid polarizing echo chambers if they know how the news they are consuming is biased. In this paper, we explore an often overlooked aspect of bias detection in documents: the semantic structure of news articles. We present DocNet, a novel, inductive, and low-resource document embedding and bias detection model that outperforms large language models. We also demonstrate that the semantic structure of news articles from opposing partisan sides, as represented in document-level graph embeddings, have significant similarities. These results can be used to advance bias detection in low-resource environments. Our code and data are made available at https://github.com/nlpresearchanon.
@@ 1 Introduction
The news media ecosystem has transformed dramatically in the past decade. Where previously, only established, professional news sources could easily disseminate information, now any outlet with an internet connection and social media presence can share their information to a broad audience Rodrigo-Ginés et al. (2023). Social media has further allowed not just the proliferation of fake news, but also the exacerbation of polarized echo chambers, where users are primarily exposed to and interact with media that already aligns with their biases Cinelli et al. (2021). As such, to maintain a well-informed society, where readers can be warned of biased reporting Spinde et al. (2022a) and can seek non-partisan or diverse reporting, it is imperative for resources for bias detection to be available.
However, given the inherently subjective nature of bias and the many different forms in which biases manifest in the media Rodrigo-Ginés et al. (2023), bias detection remains an inexact practice. Various datasets have been developed in the past five years to allow for greater research of biases at the word, sentence, article, and document level Zou et al. (2023); Wessel et al. (2023); Fan et al. (2019); Vallejo et al. (2023); Spinde et al. (2021a); Kiesel et al. (2019); Carragher et al. (2024). In addition, while automatic bias detection models historically relied on curated linguistic feature-based characterizations of text Kiesel et al. (2019); Spinde et al. (2021b), more recent studies have sought to use generalizable embeddings through deep learning-based methods, like contrastive learning Michelle YoungJin Kim (2022), pre-trained language models (e.g. BERT) Krieger et al. (2022); Vargas et al. (2023); Spinde et al. (2022b); Lei et al. (2022); Baly et al. (2020); Raza et al. (2022), LSTMs (Long Short-Term Memory neural networks) Roy and Goldwasser (2020); Lei et al. (2022), and Large Language Models (LLMs) Lin et al. (2024).
Although these deep learning methods are able to leverage word context, they do not factor in variances in writing style structures in media. As such, we propose an inductive, graph-based approach to partisan bias detection. We construct word co-occurrence graphs for each news article and find that their graph embeddings have comparable capture of bias as language model embeddings. Our detection model also statistically outperforms LLM-generated bias predictions. In this paper, our contributions are threefold:
@@ 2 Background
Bias detection is a difficult problem, not least of all because of the many different forms of bias. Recent work by Rodrigo-Ginés, et al. has identified at least 17 different categories of bias Rodrigo-Ginés et al. (2023). Bias detection in the media is complicated due to the high level of text re-use in published news articles Boumans et al. (2018). A large part of widely circulated online news consists of redistributed or edited copies from large news agencies. This text re-use and the decisions behind what to produce or what to edit often result in instances of informational and lexical biases. Informational biases occur when news articles shape how information is perceived based on what surrounding or background information is shared in the article, and lexical bias occurs when polarizing words or clauses are used Fan et al. (2019). Identification of informational bias in a high text re-use media ecosystem needs to be able to isolate these potentially minute shifts in the structure of an article from otherwise similar but differently biased text.
@@
Current news bias detection models use pre-trained word embeddings, like GloVe or Word2Vec, and language models, like BERT, to capture shifts in language in news articles. For instance, Zou et al., Zou et al. (2023) construct an article bias detection model using RoBerta-base and the POLITICS language model, which is a finetuned language model for ideology prediction Liu et al. (2022). Truica et al. combine TF-IDF (term frequency-inverse document frequency), and both non-pre-trained and pre-trained language model word embeddings to represent documents for article bias detection Truică and Apostol (2023). Meanwhile, Chen et al., use the BASIL dataset Fan et al. (2019) and predict article level biased based on the sentence level biases’ frequency, position, and sequential order Chen et al. (2020). They find that for the small BASIL dataset, lexical information (e.g. through bag of words), was insufficient to detecting article-level bias, while the features that characterize the internal structure of an article were more effective. Unfortunately, these methods require either a corpus of meticulously extracted sentence level and event features or a pre-trained language model.
@@
Article-level bias detection can also be reframed as a document classification problem. Within the more generalized document classification space, methods remain similar, with most models highly dependent on pre-trained language models or expert generated features. Doc2vec Le and Mikolov (2014) is one of the few methods that requires neither. However, transformer based approaches where word embeddings from pre-trained language models are aggregated have largely surpassed Doc2vec’s performance. Document classification researchers have begun exploring aggregating these embeddings via graph representations Wang et al. (2023); Liu et al. (2021). Much of graph based classification research has been transductive, where one graph is generated for the entire corpus of documents, connecting word nodes across documents and generating document nodes for classification Zhao and Song (2023); Dai et al. (2022); Nikolentzos et al. (2020); Yao et al. (2018); Carragher et al. (2024). Transductive models, however, are not representative of reality and are not as computationally efficient as inductive models since they require one large graph to be retrained with every addition to the corpus. InducT-GCN is an inductive classification model that creates one graph per document using document node vectors and word node vectors. Word-word edges are connected by their pointwise mutual information (PMI), word-doc edges are connected using TF-IDF, and word embeddings use pre-trained embeddings Wang et al. (2022). They train a graph convolutional network (GCN) model for text classification. However, inductive document level models struggle with longer documents due to the dense graphs that are created Piao et al. (2022). All of these models also use graphs as a method to aggregate word embeddings in a document. They do not isolate semantic structure in the constructed graphs themselves.
@@ 3 Method
In this section, we detail our methodology for detecting the biases of online news articles. The training pipeline for our supervised bias detection models has four procedures, 1) Dataset collection, 2) Graph construction, 3) Unsupervised embedding generation and 4) Supervised model training.
3.1 Datasets
Our analysis uses four datasets, three of which are centered around partisan topics: withdrawal from Afghanistan, the Oath Keepers and the military vaccine mandate, and the fourth of which is the BASIL benchmark dataset Fan et al. (2019). (For brevity, we will refer to them as the AFG, OATH, VAX, and BASIL datasets, respectively moving forward). Table 1 summarizes the four datasets.
Table 1: Summary of datasets (Notes: “Far-Left” labels do not exist in this data. “Sentences” is shortened to “sent”)
@@ 3.1 Datasets
Topic Aligned Datasets
We introduce three, new, publicly available topic-based datasets. They are collected by first pulling Tweets from 1 February 2022 until 29 December 2022 on relevant news stories based on a keyword search via the Twitter API. Table 2 displays the collection terms.
Table 2: Collection Terms for Each Dataset
Collection Terms Dataset
{afghanistan, taliban, kabul, bagram AFG
U.S. withdrawal, afghan}
{oath keepers} OATH
{vaccine, vax} VAX
We extracted the URLs from Tweets and filtered out the social media site URLs, as has been done in other works Cruickshank and Carley (2020, 2021). From there, we scraped the textual content of those websites that allowed automated scraping. This methodology results in a variety of news sources and articles that are relevant to discussions on these partisan topics Cruickshank et al. (2023). Weak article based bias labels (far left, left, left-center, center, right-center, right, far right) for each article are assigned based upon its domain bias label in lieu of human annotations Cruickshank and Carley (2021). This follows the assumption that for highly partisan topics, the level of bias of a news article and the events they discuss will be highly correlated with the partisan sentiment of its domain Zou et al. (2023). We further validate the reliabiltiy of our domain labels by comparing them with the bias levels from Adfontes, as of 10 March 2023  [^1] https://adfontesmedia.com/. Per a Pearson correlation test of domain bias labels using an ordinal encoding, our labels are highly correlated with Adfontes’ with a correlation coefficient of .79 and p-value $<e^{-80}$.
BASIL
We further validate our approach with a human annotated article level dataset, the BASIL dataset. The BASIL dataset is composed of 100 articles on the same event from three different news outlets. Each article is manually annotated with article bias labels Fan et al. (2019).
3.2 Graph Construction
We transform each news article into its own graph in order to create article embeddings after the text is cleaned and lemmatized. Stop words are also removed. Each graph is an undirected word (or lemma) co-occurrence network. With these word nodes, an edge is drawn between co-occurring words, where the weight of the edge is the positive pointwise mutual information (PPMI), Equation 1. The edge is removed if there is a non-positive PMI.
@@ 3.2 Graph Construction
We construct three variations of this graph
@@
    Node vector graph: The node attribute is the word vector per Spacy’s en_core_web_lg model  [^2] Version 3.4.1 Honnibal and Montani (2017) concatenated with the average sentiment of the given lemma across all occurrences in the article per NLTK VADER’s Sentiment Intensity Analyzer Hutto and Gilbert (2014).
3.3 Unsupervised Embeddings
Next, we generate unsupervised embeddings for each article. The embedding models are inductively trained using an 80%/20% train/test split. We do not conduct rigorous hyperparameter tuning, but rather seek to align parameters across models for better comparison.
@@ 3.3 Unsupervised Embeddings
We train three Graph2Vec models for each graph variation using the KarateClub package Rozemberczki et al. (2020). Graph2Vec conducts Weisfeiler-Lehman (WL) hashing to generate graph level features that are then passed as the “document” in Doc2Vec Le and Mikolov (2014). In the base graph’s case, the hashing step is conducted using node degrees, whereas the designated node attribute is used otherwise (e.g. the word, or the word vector from a pre-trained language model). The PPMI scores are not factored into the model as Graph2Vec does not support edge attributes. We set a dimension size of 128, 50 epochs, and 2 Weisfeiler-Lehman iterations.
@@
We also train a graph auto-encoder model (“gae”) and a variational graph auto-encoder model (“vgae”) Kipf and Welling (2016) on the word vector graph representation of the text. Both graph auto-encoders are constructed using three modules of graph convolution, graph normalization, and rectified linear unit (ReLU) layers in the encoder and an inner product decoder. In order to have a comparable embedding size and epochs to the other models, we set output feature size to be 128. We also use an Adam optimizer with a starting learning rate of .1 and a scheduler that will reduce the learning rate when the loss plateaus after a minimum of five epochs and maximum of fifty epochs.
@@
As a comparison to the graph based approaches, we also use Doc2Vec Le and Mikolov (2014), using both the distributed memory (“doc2vec_dm”) and the distributed bag of words (“doc2vec_dbow”) representations. We set a window size of 10, initial learning rate of .05, 50 epochs, minimum word count to be 3, negative sample of 5, and a vector size of 128
@@
In order to measure if graph based embeddings capture anything different from Doc2Vec, we also concatenate the separately trained Graph2Vec and Doc2Vec embeddings for a 256 dimension sized vector in the final supervised model. We combine the two variants of Doc2Vec with the three variants of Graph2Vec for a total of six different embedding models.
@@
As a final unsupervised embedding baseline comparison, we use Sentence BERT Reimers and Gurevych (2019). For each sentence in an article, we pass it through the pre-trained “All-Distilroberta-V1” encoder to map it to the default 768 dimensional vector space. A given article is subsequently represented by the average of all of its sentence embeddings.
3.4 Bias Detection
We explore the value of these various embeddings in partisan bias detection using a naive multinomial logistic regression model. We train three different inductive models using different bias label groupings 1) “Full”: the original bias categories, 2) “LCR”: Re-grouping mildly or extremely partisan articles into right or left, for three nominally encoded bias categories (right, center, left) and 3) “Binary”: Re-grouping into biased or unbiased, where left/right-center are both considered to be unbiased. After model predictions are generated, we further manually regroup the “Full” category labels into the “LCR” and “Binary” to more accurately compare accuracy between models with different numbers of categories. This allows us to explore the level of granularity in bias information in the various embeddings as well as the potential for similarities between opposing sides in partisan discussions.
@@ 3.4 Bias Detection
Due to the weak article bias labels used in the AFG, OATH, and VAX datasets, we also explore training domain level bias detection models. There is a varied number of articles between domains in the topic aligned dataset, so we aggregate the embeddings for a given domain (i.e. news source), and train a bias detector based on the combined representation of the de-duplicated domains’ article embeddings. There is only one sample of each domain after articles are aggregated. We aggregate articles in a given domain by taking the arithmetic mean of the embeddings. We also explore the effect of normalizing the embeddings by its original topic. We subtract the average embeddings within a given topic from the embedding of each article and take the average (“avgTopicDiff”) or Frobenius norm of the difference between those articles (“froNormTopicDiff”). These topic based aggregation techniques aim to remove any topical skew in the data when all three topic based datasets are trained together. With the BASIL dataset, our domain level models predict the bias level of the domain for each article without aggregating across domains since there are only three different news sources.
@@
As a final comparison, we conduct document level bias detection using pre-trained large language models finetuned for chat completion, OpenAI’s GPT Turbo 3.5 OpenAI (2021) and Meta’s Llama 3 8B Instruct Model AI@Meta (2024). We use the following zero-shot prompt template to generate bias predictions: “Please classify the following news article by its political bias. Please only classify the article as {bias labels}, and return no other text. title: {title} article: {article} bias: ”. We do not conduct domain level predictions using the large language models.
@@
We use a Zero Rule baseline for a naive model comparison where the predictions are all labeled as the majority class.

### v2

@@ Abstract
News will have biases so long as people have opinions. It is increasingly important for informed citizens to be able to identify bias as social media becomes the primary entry point for news and partisan differences increase. If people know the biases of the news they are consuming, they will be able to take action to avoid polarizing echo chambers. In this paper, we explore an often overlooked aspect of bias detection in documents: the semantic structure of news articles. We present DocNet, a novel, inductive, and low-resource document embedding and bias detection model that outperforms large language models. We also demonstrate that the semantic structure of news articles from opposing partisan sides, as represented in document-level graph embeddings, have significant similarities. These results can be used to advance bias detection in low-resource environments. Our code, data, and the corresponding datasheet are made available at:
Code and Datahttps://anonymous.4open.science/r/DocNet/
@@ 1 Introduction
The news media ecosystem has transformed dramatically in the past decade. Where previously only established, professional news sources could easily disseminate information, now any outlet with an internet connection and social media presence can share their information to a broad audience (Rodrigo-Ginés, de Albornoz, and Plaza 2023). Social media has allowed not just the proliferation of fake news, but also the exacerbation of polarizing echo chambers, where users are only exposed to and interact with media that already aligns with their biases (Cinelli et al. 2021). As such, to maintain a well-informed society, where readers can be warned of biased reporting (Spinde et al. 2022a) and can seek non-partisan or diverse reporting, it is imperative that resources for automatic article level bias detection to be available.
Given the inherently subjective nature of bias and the many different forms in which biases manifest in the media (Rodrigo-Ginés, de Albornoz, and Plaza 2023), bias detection remains an inexact practice. Various datasets have been developed in the past five years to allow for increased research of biases at the word, sentence, article, and document level (Zou et al. 2023; Wessel et al. 2023; Fan et al. 2019; Vallejo, Baldwin, and Frermann 2023; Spinde et al. 2021a; Kiesel et al. 2019; Carragher, Williams, and Carley 2024). These automatic bias detection models have historically relied on curated linguistic feature-based characterizations of text (Kiesel et al. 2019; Spinde et al. 2021b). However, more recent studies have sought to use generalizable embeddings through deep learning-based methods, like contrastive learning (Kim and Johnson 2022), pre-trained language models (Krieger et al. 2022; Vargas et al. 2023; Spinde et al. 2022b; Lei et al. 2022; Baly et al. 2020; Raza, Reji, and Ding 2022), LSTMs (Long Short-Term Memory neural networks) (Roy and Goldwasser 2020; Lei et al. 2022), and Large Language Models (LLMs) (Lin et al. 2024).
Although these deep learning methods leverage word context, they do not explicitly factor in variances in writing style structures in media. As such, we propose an inductive, graph-based approach to partisan bias detection. We construct word co-occurrence graphs for each news article and discover that their graph embeddings have comparable performance in capturing partisan bias as language model embeddings. Our detection model also statistically outperforms LLM-generated bias predictions. In this paper, our contributions are threefold:
@@ 2 Background
Bias detection is a difficult problem, not least of all because of the many different forms of bias. Recent work by Rodrigo-Ginés, et al. has identified at least 17 different categories of bias (Rodrigo-Ginés, de Albornoz, and Plaza 2023). Bias detection in the media is further complicated due to the high level of text re-use in published news articles (Boumans et al. 2018) and the different frames available to discuss an issue (Card et al. 2015). A large part of widely circulated online news consists of redistributed or edited copies from large news agencies. This text re-use and the decisions behind what to publish or how to edit an article result in instances of informational and lexical biases. Informational biases occur when news articles change the surrounding or background information in order to shape how information is framed, and thereby perceived. Lexical bias occurs when polarizing words or clauses are used to illicit specific emotion towards an issue (Fan et al. 2019). Partisan biases in news articles are generated through these shifts in content and languages. Identification of these rhetorical methods in a high text re-use media ecosystem requires isolating minute shifts in the language and structure of otherwise similar articles.
@@
Current news bias detection models use pre-trained word embeddings, like GloVe or Word2Vec, and language models, like BERT, to capture biased language in news articles. For instance, Zou et al., (Zou et al. 2023) construct an article bias detection model using RoBerta-base and the POLITICS language model, which is a finetuned language model for ideology prediction (Liu et al. 2022). Truica et al. combine TF-IDF (term frequency-inverse document frequency), and both non-pre-trained and pre-trained language model word embeddings to represent documents for article bias detection (Truică and Apostol 2023). Meanwhile, Chen et al., use the BASIL dataset (Fan et al. 2019) and predict article level biases using the sentence level biases’ frequency, position, and sequential order (Chen et al. 2020). They find that for the small BASIL dataset, lexical information (e.g. through bag of words), was insufficient in detecting article-level bias, while the features that characterize the internal structure of an article were more effective. Yet, all of these methods hinge on either a corpus of meticulously extracted sentence level and event features or a suitable pre-trained language model.
@@
Article-level bias detection can also be reframed as a document classification problem. Within the more generalized document classification space, methods remain similar, with most models highly dependent on pre-trained language models, like Longformers (Beltagy, Peters, and Cohan 2020) or expert generated features. Doc2vec (Le and Mikolov 2014) is one of the few methods that requires neither. However, transformer based approaches, where word embeddings from pre-trained language models are aggregated, have largely surpassed Doc2vec’s performance.
Document classification researchers have also begun aggregating these embeddings via graph representations (Wang, Ding, and Han 2023; Liu et al. 2021). Much of graph based document classification research has been transductive, where one graph is generated for the entire corpus of documents. Where word nodes are connected across documents and document nodes are generated for classification (Zhao and Song 2023; Dai et al. 2022; Nikolentzos, Tixier, and Vazirgiannis 2020; Yao, Mao, and Luo 2019; Carragher, Williams, and Carley 2024). Transductive models, however, are not representative of reality and are not as computationally efficient as inductive models since they require one large graph to be retrained with every addition to the corpus.
InducT-GCN is one of the few inductive document classification models discussed in literature (Wang, Han, and Poon 2022). It creates one graph for the entire training corpus using document node vectors and word node vectors. Word-word edges are connected by their pointwise mutual information (PMI) and word-doc edges are connected using TF-IDF. They train a graph convolutional network (GCN) model to learn the word embeddings and conduct classification. These word embeddings and learned weights are used to generate predictions on batches of test documents. However, other researches have found that inductive document level models struggle with longer documents due to the dense graphs that are created (Piao et al. 2022). This is apparent in the case of Induct-GCN, where while still outperforming baselines, there is observably worse performance on longer text document datasets like 20 News Group and Ohsumed, than for shorter ones like R52 (Wang, Han, and Poon 2022). The aforementioned graph based models have also only used graphs as a method to aggregate word embeddings in a document. They do not isolate the semantic structure of individual documents in the constructed graphs.
@@ 3 Method
In the following section, we detail our methodology for detecting the biases of online news articles. The pipeline for our inductive supervised bias detection models has four procedures, 1) Dataset collection, 2) Graph construction, 3) Unsupervised embedding generation and 4) Supervised model training.
Datasets
Our analysis uses four datasets, three of which are centered around partisan topics: withdrawal from Afghanistan, the Oath Keepers and the military vaccine mandate, and the fourth of which is the BASIL benchmark dataset (Fan et al. 2019). (For brevity we will refer to them as the AFG, OATH, VAX, and BASIL datasets, respectively, moving forward).
Topic Aligned Datasets
We introduce three publicly available topic-based datasets, the AFG, Oath, and VAX news article datasets. We chose these three topics for experimentation due to the significant attention, controversy, and corresponding media coverage surrounding each. Each topic’s respective articles are collected by first pulling Tweets from 1 February 2022 until 29 December 2022 on relevant news stories based on a keyword search via the Twitter API. Table 1 displays the collection terms.
Collection Terms Dataset
{afghanistan, taliban, kabul, bagram AFG
U.S. withdrawal, afghan}
{oath keepers} OATH
{vaccine, vax} VAX
Table 1: Collection Terms for Each Dataset
We extracted the URLs from Tweets and filtered out the social media site URLs, as has been done in other works (Cruickshank and Carley 2020, 2021). From there, we scraped the textual content of those websites that allowed automated scraping. This methodology results in a variety of news sources and articles that are relevant to discussions on these partisan topics (Cruickshank, Zhu, and Bastian 2023). We apply domain bias labels from a dataset compiled by Cruickshank et al (Cruickshank and Carley 2021) and available under CC by 4.0 licensing. Domains are the publishing sites that news articles are scraped from. We use the term news domain in lieu of news sources, as articles are often shared from distributing news agencies, like the Associated Press, to multiple sites (i.e. domains).
A weak bias label (far left, left, left-center, center, right-center, right, far right) for each article is assigned based upon its domain’s bias label. Our usage of domain labels as weak labels for article bias follows the assumption that for highly partisan topics, the bias of a news article and the events that are discussed are highly correlated with the perspective of its publisher, or domain (Zou et al. 2023). We further validate the reliability of Curickshank et al’s domain bias labels by comparing them with the domain bias levels published on Adfontes, as of 10 March 2023  [^1] https://adfontesmedia.com/. Per a Pearson correlation test of news domain bias labels using an ordinal encoding, we validate that our labels are highly correlated with Adfontes’ with a correlation coefficient of .79 and p-value $<e^{-80}$. This is in spite of the three year difference in times at which these domain bias labels were collected (2020 vs 2023) and the likely drift in media bias over time. This test increases confidence that the domain bias labels we use have minimal bias themselves, as both data sources report similar labels.
BASIL
We further validate our approach with a dataset that has human annotated article level biases, the BASIL dataset (Fan et al. 2019). The BASIL dataset is composed of 300 articles from 2010 to 2019 that cover 100 events, with one article each from Fox News, New York Times, and Huffington Post. Each article is human annotated with article and domain bias labels. Key characteristics of these four datasets are summarized in table 2.
@@
Table 2: Summary of datasets (Note: “Far-Left” labels do not exist in this data. “Sentences” is shortened to “sent”)
Graph Construction
Next, we transform each news article into its own graph after the text is cleaned and lemmatized in order to capture each article’s individual semantic structure in the article’s embedding. Each graph is an undirected word (or lemma) co-occurrence network. With these word nodes, an edge is drawn between co-occurring words, where the weight of the edge is the positive pointwise mutual information (PPMI), Equation 1. The edge is removed if there is a non-positive PMI.
@@
We reserve the cleaned text of the article for comparison model experimentation and construct three variations of the article graph .
@@
    Node vector graph: The node attribute is the word vector per Spacy’s en_core_web_lg model  [^2] Version 3.4.1 (Honnibal and Montani 2017) and concatenated with the average sentiment of the given lemma across all occurrences in the article per NLTK VADER’s Sentiment Intensity Analyzer (Hutto and Gilbert 2014).
Unsupervised Embeddings
From the preprocessed texts, we generate unsupervised embeddings of each article. The embedding models are inductively trained using an 80%/20% train/test split. We do not conduct hyperparameter tuning. Instead, we seek to generally align parameters, like epochs and dimension size, across models for better comparison, or otherwise use default package hyperparameters.
@@
We train three Graph2Vec models for each graph variation using the KarateClub package (Rozemberczki, Kiss, and Sarkar 2020). Graph2Vec conducts Weisfeiler-Lehman (WL) hashing to generate graph level features that are then passed as the “document” in Doc2Vec (Le and Mikolov 2014). In the base graph’s case, the hashing step is conducted using node degrees, whereas the designated node attribute is used otherwise (e.g. the word, or the word vector from a pre-trained language model). The PPMI scores are not factored into the model as Graph2Vec does not currently support edge attributes. We set a dimension size of 128, 50 epochs, and 2 Weisfeiler-Lehman iterations.
@@
We also train a graph auto-encoder model (“gae”) and a variational graph auto-encoder model (“vgae”) (Kipf and Welling 2016) on the node vector graph representation of the text, both with and without the sentiment vector (“sent”). Both graph auto-encoders are constructed using three modules of graph convolution, graph normalization, and rectified linear unit (ReLU) layers in the encoder, and an inner product decoder. In order to have a comparable embedding size and epochs to the other models, we set output feature size to be 128. We use an Adam optimizer with a starting learning rate of .1 and a scheduler that reduces the learning rate when the loss plateaus after a minimum of five epochs and maximum of fifty epochs.
@@
As a comparison to the graph based approaches, we use Doc2Vec (Le and Mikolov 2014), using both the distributed memory (“doc2vec_dm”) and the distributed bag of words (“doc2vec_dbow”) representations. We set a window size of 10, initial learning rate of .05, 50 epochs, minimum word count to be 3, negative sample of 5, and a vector size of 128.
@@
In order to measure if graph based embeddings capture anything different from Doc2Vec, we also concatenate the separately trained Graph2Vec and Doc2Vec embeddings into a 256 dimension sized vector in the final supervised model. We combine the two variants of Doc2Vec with the three variants of Graph2Vec for a total of six different embedding models.
@@
As a final unsupervised embedding comparison, we use Sentence BERT (Reimers and Gurevych 2019). For each sentence in an article, we pass it through the pre-trained “All-Distilroberta-V1” encoder to map it to a default 768 dimensional vector space. A given article is subsequently represented by the average of all of its sentence embeddings.
Bias Detection
We explore the value of these various embeddings in partisan bias detection using a naive multinomial logistic regression model. We train three inductive models using varied bias label groupings 1) “Full”: the original bias categories, 2) “LCR”: Re-grouping mildly or extremely partisan articles into right or left, for three nominally encoded bias categories (right, center, left) and 3) “Binary”: Re-grouping into biased or unbiased, where left/right-center are both considered to be unbiased. After predictions are generated, we further manually regroup the “Full” category labels into the “LCR” and “Binary” to more accurately compare accuracy of models with different numbers of categories. These variations allow us to explore the level of granularity of the bias information captured in the various embeddings. It also allows us to interrogate the potential for similarities between opposing viewpoints in partisan discussions.
@@
We also explore training domain level bias detection models due to the weak article bias labels used in the AFG, OATH, and VAX datasets. In the topic aligned datasets, there is a mixed number of articles between domains. As a result, we develop various methods to deduplicate domains for training a news domain bias detector by aggregating the article embeddings of each domain. In other words, the domain bias detectors are trained on a combined representation of each domain’s article embeddings.
We apply three methods of aggregating articles for each domain. First, we aggregate articles in a given domain by taking the arithmetic mean of the embeddings (“arithmeticMean”, equation 2). We also explore the effect of normalizing the embeddings by its original topic. We subtract the average embeddings within a given topic from the embedding of each article and take the average (“avgTopicDiff”, equation 4), and for another method, take the Frobenius norm (“froNormTopicDiff”, equation 5). The latter two topic based aggregation techniques aim to remove any topic specific skew in the data when all three topic aligned datasets are aggregated.
arithmeticMean: $\begin{split}\text{arithmeticMean: }d_{j}=\frac{1}{n_{j}}\sum_{i=1}^{n_{j}}\sum_{t=1}^{n_{t}}x_{ijt}\\ \end{split}$ (2)
$\begin{split}T_{t}=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}x_{ijt}\\ \hat{x}_{ijt}=x_{ijt}-T_{t}\end{split}$ (3)
avgTopicDiff: $\begin{split}\text{avgTopicDiff: }d_{j}=\frac{1}{n_{j}}\sum_{i=1}^{n_{j}}\sum_{t=1}^{n_{t}}\hat{x}_{ijt}\\ \end{split}$ (4)
froNormTopicDiff: $\begin{split}\text{froNormTopicDiff: }d_{j}=\sqrt{\sum_{i=1}^{n_{j}}|\hat{x}_{ijt}|^{2}}\\ \end{split}$ (5)
where,
$d_{j}=\text{embedding of domain, $j$}$
$x_{ijt}=\text{embedding of article $i$ from domain $j$ and topic$t$}$
$n=\text{number of articles}$
$T_{t}=\text{average topic embedding of topic $t$}$
With the BASIL dataset, our domain level models predict the bias level of the domain for each article without aggregating embeddings as there are only three different news domains. This increases the risk of the model capturing spurious characteristics, but in reviewing our cleaning procedure, we believe the risk of domain specific text cues being present is sufficiently mitigated.
@@
As a final comparison, we conduct document level bias detection using pre-trained large language models finetuned for chat completion. We experiment with both OpenAI’s GPT Turbo 3.5 (OpenAI 2021) and Meta’s Llama 3 8B Instruct Model (AI@Meta 2024). We pass the following zero-shot prompt template to generate bias predictions: “Please classify the following news article by its political bias. Please only classify the article as {bias labels}, and return no other text. title: {title} article: {article} bias: ”. We do not conduct domain level predictions using the large language models.
@@
We use a Zero Rule baseline for a naive model standard, where the test set predictions are all labeled as the majority class in the training set.